You are standing at a distance of 10.0 m from a speaker. The sound intensity is some value. You walk to a distance 4 times farther. How is the sound intensity affected?
1
Expert's answer
2019-06-19T10:25:42-0400

In accordance with the law of conservation of energy

"I\cdot4\pi R^{2} = const,"

respectively

"I_2 = I_1(\frac{R_1}{R_2})^{2} = \frac{I_1}{16},"

i.e. intensity will decrease by 16 times.